OpenStudy (anonymous):

Simplify the given expression. square root of negative 54 end root plus square root of negative 96

OpenStudy (anonymous):

mathematical one..

OpenStudy (anonymous):

\[\sqrt{-54} + \sqrt{-96}\]

OpenStudy (anonymous):

i*sqrt(54) + i*sqrt(96) = i*9 sqrt(6) + 4i*sqrt(6) =13i sqrt(6)

OpenStudy (anonymous):

\[\left(\sqrt{-54}+\sqrt{-96}\right)^2=\sqrt{-294}=7 i \sqrt{6} \]

OpenStudy (anonymous):

\[5i \sqrt{6}\] \[-5\sqrt{6}\] \[-7\sqrt{6}\] \[7i \sqrt{6}\]

OpenStudy (anonymous):

what does the i mean?

OpenStudy (anonymous):

@robtobey - i didnt understand..

OpenStudy (anonymous):

why did you "square" it???

OpenStudy (anonymous):

i stand for imaginary num i=-1

OpenStudy (anonymous):

i=sqrt(-1)

OpenStudy (anonymous):

Square the problem expression -> -294 Square root of -294 = 7 i radical 6 i is the imaginary number equivalent to radical 1.

OpenStudy (anonymous):

good way @rob

OpenStudy (anonymous):

i is an imaginary num equals to /sqrt{-1}

OpenStudy (anonymous):

Mistake, "to radical -1.' is what was intended.

OpenStudy (anonymous):

why you took square of whole ?

OpenStudy (anonymous):

makes the sum easier @deadman340 he did it in less no. of steps hence consuming less time

OpenStudy (anonymous):

so by taking square it will not effect the anss?